OpenStudy (anonymous):

can someone show me how to do this problem. square root of 6a - 4 square root of 54a - 4 square root of 216a

OpenStudy (ahsome):

So:\[\sqrt{6a-4}, \sqrt{54a-4}, \sqrt{216a}\] Right?

OpenStudy (ahsome):

Is that the question? Are you trying to simplify it?

OpenStudy (anonymous):

yes

OpenStudy (ahsome):

1) \[\sqrt{6a-4}\] Split into this: \[\sqrt{6a}*\sqrt{-1}*\sqrt{4}\] Simplify that to: \[\sqrt{6a}*2*i\] This equals to: \[2i\sqrt{6a}\]

OpenStudy (ahsome):

You would do the same thing with the other 2, with \[\sqrt{216a}\] being the easier, as ther is no \[i\]

OpenStudy (ahsome):

Wait, I think that might not be right ;) You have this equation:\[\sqrt{6a-4}\] Split that into \[\sqrt{6a}+\sqrt{-4}\] Change that to \[\sqrt{6a}+\sqrt{4}\sqrt{-1}\] This can be turned into \[\sqrt{6a}+2i\]

OpenStudy (ahsome):

I am new to this, so someone more qualified, please help

OpenStudy (anonymous):

Is the question: \[\sqrt{6a-4}\sqrt{54a-4}\sqrt{216a}\] And you have to simplify that?

OpenStudy (ahsome):

Hmm, that's possible. You would have to simplify everything first as much as possbile before doing it to make it easier

OpenStudy (ahsome):

I know that:\[\sqrt{6a-4}\] is equal to: \[\sqrt{2}\sqrt{3a-2}\]

OpenStudy (anonymous):

That is a correct operation (above). You cannot do this: \[\sqrt{6a-4}\neq \sqrt{6a}-\sqrt{4}\] So your answers with imaginary components are invalid.

OpenStudy (anonymous):

Which is why the second interpretation of the question is more likely.

OpenStudy (ahsome):

I was thinking that. Sorry ;)

OpenStudy (ahsome):

That would be one long and complicated question, I think he is just missing some important piece of info on the question.

OpenStudy (anonymous):

Haha, that's all right. If you are not sure of an operation then simply type it into wolfram alpha like this: http://www.wolframalpha.com/input/?i=does+sqrt%286a-4%29%3Dsqrt%286a%29-sqrt%284%29

OpenStudy (anonymous):

And yeah, that is a good point, I suppose we should just see if s/he comes back

OpenStudy (ahsome):

WolframAlpha & MathMatica is an AWESOME combination. Great thing that you can use WolframAlpha inside MathMatica